# An entanglement witness and an effective criterion for bipartite states of different dimensions

### v4

@@ 1 Construction of positive maps for arbitrary different dimensions
where $|\beta\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d_{1}-1}|\mu,\mu\rangle$ is a maximally entangled state. To do this calculation, we extend the orthonormal sets $\{\Gamma_{i}\}$ and $\{\Omega_{k}\}$ respectively to orthonormal bases for the space of matrices $M_{d_{1}}$ and $M_{d_{2}}$, namely to $\{\Gamma_{\mu}\}=\{\Gamma_{0}=\frac{I}{\sqrt{d_{1}}},\Gamma_{i}\}$ and $\{\Omega_{\mu}\}=\{\Omega_{0}=\frac{I}{\sqrt{d_{2}}},\Omega_{i}\}$ and taking remark 1 into account, we write
$W$ $=$ $\frac{1}{d_{1}}\sum_{\mu,\nu}|\mu\rangle\langle\nu|\otimes\Phi(|\mu\rangle\langle\nu|)$ (10)
$=$ $\frac{1}{d_{1}}\sum_{\mu,\nu}|\mu\rangle\langle\nu|\otimes\sum_{\alpha=0}^{d_{1}^{2}-1}\sum_{\beta=0}^{d_{2}^{2}-1}\Lambda_{\beta\alpha}\operatorname{Tr}(|\mu\rangle\langle\nu|\Gamma_{\beta})\Omega_{\alpha}$
$=$ $\frac{1}{d_{1}}\sum_{\alpha,\beta}\Lambda_{\beta\alpha}\Gamma^{T}_{\beta}\otimes\Omega_{\alpha}.$ (11)
@@
where $\Lambda_{00}=\frac{1}{d_{2}},\Lambda_{0i}=0,\Lambda_{i0}=0$. After multiplication by an overall factor $d_{1}d_{2}$, this leads to the following witness
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}{\Lambda}_{ik}\Gamma^{T}_{i}\otimes\Omega_{k}.$
Since the set $\Gamma^{T}_{i}$ is also an orthonormal set, we can rename it to $\Gamma_{i}$. Thus, the following matrix can be a entanglement witness too:
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}{\Lambda}_{ik}\Gamma_{i}\otimes\Omega_{k}.$ (12)
@@
Therefore for any two dimensions, any operator of the above form can be an entanglement witness, provided that the condition (9) holds. That condition in fact proves Block positivity of this Witness. Whether or not a $W$ constructed in this way, has also a negative eigenvalue depends on the matrix $\Lambda$ that we choose. In all previous papers that were mentioned in the introduction, $\Lambda$ should be orthogonal. These comprise a much smaller class compared to the $\Lambda$ matrices in this paper. Moreover, finding orthogonal matrices in high dimensions is not a straightforward task. In this paper, one can use any matrix with any properties to construct entanglement witnesses, as long as the largest eigenvalue of $\Lambda\Lambda^{T}$ satisfies eq. (9).
In the next section, we derive the same result of eq. (12) in an alternative way by focusing on the EW itself, rather than the positive map which leads to it via its Choi-matrix. In view of this alternative construction, one can first choose the local operators (according to the experimental restrictions) and then tuning the coefficients for defining the final entanglement witness. This is a significant advantage compared to methods where a positive map may lead to a witness which requires non-feasible measurements.
2 An alternative construction of Entanglement Witnesses
In this section we directly apply Mehta’s Lemma to a form of EW which we take as follows:
$W=I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}\Lambda_{ij}\Gamma_{i}\otimes\Omega_{j}.$
We demand that $W$ be block-positive. This means that for any vector $|v\rangle\in H_{d_{2}}$, the following matrix should be positive
$B\equiv\langle v|W|v\rangle=I_{d_{1}}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}\Lambda_{ij}\Gamma_{i}\langle v|\Omega_{j}|v\rangle.$
Now we use the following:
$\operatorname{Tr}(B)=d_{1},$
and
$\operatorname{Tr}(B^{2})=d_{1}+\sum_{i,j=1}^{d_{2}^{2}-1}(\Lambda^{T}\Lambda)_{ij}\langle v|\Omega_{i}|v\rangle\langle v|\Omega_{j}|v\rangle,$
where we have used the relations $\operatorname{Tr}(\Gamma_{i})=0,\ \ \ \operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij}$. We now resort to the Mehta’s Lemma and demand that the following condition hold for all the vectors $|v\rangle$
$\operatorname{Tr}(B^{2})\leq\frac{\operatorname{Tr}(B)^{2}}{d_{1}-1}.$
This leads to
[…]
Defining the vector ${\bf q}$ with components $q_{i}:=\langle v|\Omega_{i}|v\rangle$, this can be written as
[…]
or in a stronger form
$d_{1}+\lambda_{max}(\Lambda^{T}\Lambda){\bf q}.{\bf q}\leq\frac{d_{1}^{2}}{d_{1}-1}.$ (13)
It remains to calculate ${\bf q}.{\bf q}$, and this is easily found by noting that
${\bf q}.{\bf q}=\sum_{i=1}^{d_{2}^{2}-1}\langle v|\Omega_{i}|v\rangle\langle v|\Omega_{i}|v\rangle=\sum_{i=1}^{d_{2}^{2}-1}\langle v,v|\Omega_{i}\otimes\Omega_{i}|v,v\rangle.$
Using the identity
$\frac{1}{d}I+\sum_{i}\Omega_{i}\otimes\Omega_{i}=P,$ (14)
where $P$ is the permutation operator, we find that
${\bf q}.{\bf q}=\langle v,v|P-\frac{1}{d_{2}}I|v,v\rangle=1-\frac{1}{d_{2}}.$
Inserting this in (13), leads to the final form for the upper bound of $\lambda_{max}$, namely
$\lambda_{max}(\Lambda^{T}\Lambda)\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)},$ (15)
which is identical to the bound obtained by the positive map in section 1. However we should also ensure that $W$ is not trivially positive, that is we should remove the possibility that all its eigenvalues be positive. We again use Mehta’s lemma. To this end we use
$\operatorname{Tr}(W)=d_{1}d_{2},$
and
$\operatorname{Tr}(W^{2})=d_{1}d_{2}+\operatorname{Tr}(\Lambda^{T}\Lambda),$
and ensure that the following inequality be satisfied
$\operatorname{Tr}(W^{2})>\frac{\operatorname{Tr}(W)^{2}}{d_{1}d_{2}-1},$
or
$\frac{d_{1}d_{2}}{d_{1}d_{2}-1}<\operatorname{Tr}(\Lambda^{T}\Lambda).$ (16)
Combining (15) and (16), we arrive at the final condition for the parameters of the matrix $\Lambda$, in order that $W$ can be an entanglement witness:
$\frac{d_{1}d_{2}}{d_{1}d_{2}-1}<\operatorname{Tr}(\Lambda^{T}\Lambda),\ \ \ \ {\rm and}\ \ \ \lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}.$ (17)
Note that the second condition is sufficient for the block-positivity of $W$, while the first condition is necessary for $W$, ensuring that it is not a positive matrix.
3 A simple criterion
Consider now the form of the witness in equation (12) subject to the condition
$\lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)},$ (18)
which is sufficient for its block-positivity. Let an arbitrary bi-partite state be expressed as follows in an orthogonal trace-less basis:
$\rho=\frac{1}{d_{1}d_{2}}I_{d_{1}}\otimes I_{d_{2}}+\sum_{i=1}^{d_{1}^{2}-1}{a}_{i}\Gamma_{i}\otimes I_{d_{2}}+\sum_{j=1}^{d_{2}^{2}-1}{b}_{j}I_{d_{1}}\otimes\Omega_{j}+\sum_{i=1}^{d_{1}^{2}-1}\sum_{j=1}^{d_{2}^{2}-1}{r}_{ij}\Gamma_{i}\otimes\Omega_{j}.$
It is readily seen that
$\operatorname{Tr}(W\rho)=1+\sum_{ij}\Lambda_{ij}r_{ij}=1+\operatorname{Tr}(R\Lambda^{T}),$ (19)
where the matrix $R$ is composed of $r_{ij}$ elements. In order to make $\operatorname{Tr}(W\rho)<0$, which at the same time proves that $W$ is indeed an entanglement witness (i.e. has a negative eigenvalue) and detects the entanglement of $\rho$, we try to choose $\Lambda$ in such a way to minimize $tr(R\Lambda^{T})$, while maintaining the condition (18). To achieve this, we first express the matrix $R$ using singular value decomposition:
$R=USV.$
Here we have used the fact that $R$ is a real matrix, which implies the matrices $U\in M_{d_{1}\times d_{1}}$ and $V\in M_{d_{2}\times d_{2}}$ are real orthogonal matrices. The matrix $S\in M_{d_{1}\times d_{2}}$ is a rectangular diagonal matrix with non-negative entries on the main diagonal. Now we choose the matrix $\Lambda^{T}$ as follows:
$\Lambda^{T}=V^{T}LU^{T},$
where $L\in M_{d_{2}\times d_{1}}$ is chosen to be a rectangular diagonal matrix whose entries on its main diagonal are equal to $l_{i}=-\sqrt{\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}}$. This leads to
$\Lambda^{T}\Lambda=V^{T}(LL^{T})V,$
where $LL^{T}\in M_{d_{2}\times d_{2}}$ is a matrix whose diagonal elements are either zero or equal to $\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}$. This means that the condition (9) on the maximum eigenvalue of $\Lambda^{T}\Lambda$ is satisfied. We then find that
$\operatorname{Tr}(\Lambda^{T}R)=\operatorname{Tr}(SL)=-\sqrt{\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}}\operatorname{Tr}(S).$
In this way we have obtained the lowest value of $\operatorname{Tr}(\Lambda^{T}R)$, while maintaining the condition (18) of block-positivity of $W$. Therefore, if this value is less than $-1$, which in view of (19) is equivalent to $\operatorname{Tr}(W\rho)<0$, we conclude that the state $\rho$ is entangled. In addition, one can show that $Tr(S)=Tr(\sqrt{RR^{T}})$. Therefore, we have the following simple criterion:
${\rm If\ \ }\gamma:=\operatorname{Tr}(S)>\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}},\ \ \ {\rm the\ state\ is\ entangled.}$ (20)
The condition that we derived does not require finding the elements of the matrix $\Lambda$, and only requires recasting the density matrix in terms of a basis chosen by orthogonal trace-less matrices, and calculating the singular values of the matrix $R$.
It should be mentioned that the singular values of $R$ and hence the value of $\gamma$ do not depend on the bases which we choose for the two parts for expanding the density matrix $\rho$, as long as these bases are orthonormal and traceless.
Remark: Interestingly, this criterion is a special case of Theorem 1 in [22], where the authors show that this is a powerful criterion and unifies several criteria known before.
@@ 3 A simple criterion
We will now show by several examples that this criterion, originated from the EW defined in (12), is indeed powerful enough to detect entanglement of a class of PPT entangled states which has not been previously detected.
@@ 4 Examples
In this section, we consider a few examples and show the power of our criterion. In all the following examples, except example 4.2, the states are PPT entangled states reported in previous works. We will show that the criterion given in (20), can detect entanglement in a larger interval of the parameters previously reported, and it also can detect PPT entangled states with different dimensions, i.e. $d_{1}\neq d_{2}$. For checking the following examples, the reader should note that we have used the following order for Gellmann matrices:
@@ 4.1 Example 1:
In this example, we demonstrate that the method explained in this paper can identify PPT entangled states, and thus, it can generate indecomposable EWs. It has been shown in [23] that the following state is PPT entangled state when $x\neq 1$ (for $x=1$ this is a separable state):
$\rho=\frac{1}{N}\left(\begin{array}[]{ccc|ccc|ccc}1&.&.&.&1&.&.&.&1\\ .&\frac{1}{x}&.&.&.&.&.&.&.\\ .&.&x&.&.&.&.&.&.\\ \hline\cr.&.&.&x&.&.&.&.&.\\ 1&.&.&.&1&.&.&.&1\\ .&.&.&.&.&\frac{1}{x}&.&.&.\\ \hline\cr.&.&.&.&.&.&\frac{1}{x}&.&.\\ .&.&.&.&.&.&.&x&.\\ 1&.&.&.&1&.&.&.&1\\ \end{array}\right).$ (23)
Here dots represent zeros, $N=3+x+x^{-1}$ is the normalization factor and $x>0$. The above state can be written in the basis of normalized Gell-Mann matrices as follows:
[…]
$+$ $\left(1-\frac{\xi}{2}\right)(\Gamma_{7}\otimes\Gamma_{7}+\Gamma_{8}\otimes\Gamma_{8})-\frac{\sqrt{3}}{2}\eta(\Gamma_{7}\otimes\Gamma_{8}-\Gamma_{8}\otimes\Gamma_{7})\Big{]},$
where $\xi=x+x^{-1}$ and $\eta=x-x^{-1}$. The eigenvalues of the matrix $R$ are easily found to be
$\lambda(R)=\frac{1}{N}\{1(3),-1(3),(1-\frac{\xi}{2})\pm i\frac{\sqrt{3}}{2}\eta\},$
where the numbers in parentheses indicate the multiplicity of each singular value. Thus, taking into account the normalization factor $N$, we find
$Tr(S)=\frac{6+2\sqrt{(1-\frac{\xi}{2})^{2}+\frac{3}{4}\eta^{2}}}{3(1+\xi)}.$
Condition (20) for entanglement of the state now amounts to
$\frac{1}{4}(\xi^{2}-\eta^{2})<\xi-1,$
or after simplification (inserting $\xi=x+x^{-1}$ and $\eta=x-x^{{}_{1}}$) leads to the following condition for the state (23) to be entangled:
$2<x+x^{-1}.$
This shows that the state $\rho_{x}$ is entangled for all values of $x$, except for $x=1$, which coincides with the result of [23].
4.2 Example 2:
@@
To calculate $\gamma$ in equation (20), we need to obtain the coefficients $r_{ij}$:
@@
$\gamma=\operatorname{Tr}(S)=\frac{p}{d}(d^{2}-1),$
which, according to eq (20), should be greater than $\frac{d-1}{d}$ so the state is entangled. Thus the range of $p$ for the state to be entangled is $p>\frac{1}{d+1}$. This is the same value obtained from the Peres criterion which in this case follows from
@@
4.3 Example 3
In this example, we provide a PPT entangled state which can be detected by the criterion in equation (20). To construct a PPT entangled state we define the following Unextendible Product Basis in $C_{3}\otimes C_{4}:$
$|\phi_{1}\rangle=\frac{1}{\sqrt{2}}|0\rangle|1-2\rangle,\quad|\phi_{2}\rangle=\frac{1}{\sqrt{6}}|0\rangle(2|0\rangle-|1+2\rangle),\quad|\phi_{3}\rangle=\frac{1}{\sqrt{2}}|1\rangle|1-2\rangle,\quad$
$|\phi_{4}\rangle=\frac{1}{\sqrt{2}}|2\rangle|1-2\rangle,\quad|\phi_{5}\rangle=\frac{1}{\sqrt{6}}|2\rangle(|1+2\rangle-2|3\rangle),\quad|\phi_{6}\rangle=\frac{1}{\sqrt{2}}|1-2\rangle|0\rangle,\quad$ (25)
$|\phi_{7}\rangle=\frac{1}{\sqrt{2}}|0-1\rangle|3\rangle,\quad|\phi_{8}\rangle=\frac{1}{\sqrt{12}}|0+1+2\rangle|0+1+2+3\rangle,$ (26)
and construct the PPT entangled state $\rho=I-\sum_{i}|\phi_{i}\rangle\langle\phi_{i}|$ [8], whose explicit form is given by
$\rho=\frac{1}{48}\left(\begin{array}[]{c|c|c|c}3\ \epsilon&-\epsilon&-\epsilon&-\epsilon\\ \hline\cr-\epsilon&D&D^{\prime}&-\epsilon\\ \hline\cr-\epsilon&D^{\prime T}&D&-\epsilon\\ \hline\cr-\epsilon&-\epsilon&-\epsilon&3\ \epsilon\\ \end{array}\right),$
where $\epsilon$ is the 3 dimensional square matrix all of its elements are equal to 1, and
$D=\left(\begin{array}[]{ccc}5&-1&-1\\ -1&5&-1\\ -1&-1&5\end{array}\right),\quad D^{\prime}=\left(\begin{array}[]{ccc}-1&5&-1\\ -1&-1&5\\ 5&-1&-1\end{array}\right).$
The matrix $R$ is simply obtained by noting that its elements are given by $r_{mn}=\operatorname{Tr}(\rho(\Gamma_{m}\otimes\Omega_{n}))=-\sum_{i}\langle\phi_{i}|\Gamma_{m}\otimes\Omega_{n}|\phi_{i}\rangle$. The result is
$R=\frac{1}{24}\left(\begin{array}[]{c|c|c}-\epsilon_{3\times 6}&0_{3\times 6}&A_{3\times 3}\\ \hline\cr 0_{3\times 6}&0_{3\times 6}&0_{3\times 3}\\ \hline\cr B_{2\times 6}&0_{2\times 6}&C_{2\times 3}\end{array}\right),$
where $\epsilon_{3\times 6}$ is a $3\times 6$ matrix all of whose elements are 1, and similarly ${0}_{3\times 6}$ is a matrix all of whose elements are zero. Also
$A$ $=$ $\left(\begin{array}[]{ccc}0&0&-3\sqrt{\frac{3}{2}}\\ 0&0&0\\ 3&\sqrt{3}&\sqrt{\frac{3}{2}}\end{array}\right),$ (30)
$B$ $=$ $\left(\begin{array}[]{cccccc}{2}&{2}&-1&0&0&0\\ \frac{2}{\sqrt{3}}&\frac{2}{\sqrt{3}}&\frac{1}{\sqrt{3}}&0&\frac{-4}{\sqrt{3}}&\frac{-4}{\sqrt{3}}\end{array}\right),$ (33)
$C$ $=$ $\left(\begin{array}[]{ccc}0&0&-\frac{1}{2}\sqrt{\frac{3}{2}}\\ -\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{5}{6\sqrt{2}}\end{array}\right).$ (36)
The singular values of this matrix are obtained numerically and one finds that
$\gamma=\operatorname{Tr}(S)=1.429>\sqrt{\frac{3\times 4}{2\times 3}}=\sqrt{2},$
which shows that while $\rho$ is a PPT entangled state, it can be detected by the criterion (20).

### v5

@@ 4 Entanglement Witnesses from positive maps
where $|\beta\rangle=\frac{1}{\sqrt{d_{1}}}\sum_{i=0}^{d_{1}-1}|i,i\rangle$ is a maximally entangled state. Direct calculation, after multiplication by an overall factor $d_{1}d_{2}$, leads to the following witness
$W=I_{d_{1}}\otimes I_{d_{2}}+t_{\mu}I_{d_{1}}\otimes\Omega_{\mu}+d_{1}{\Lambda}_{\mu i}\Gamma_{i}\otimes\Omega_{\mu}.$ (24)
In the last step, we have used the freedom to replace the matrices $\Gamma_{i}^{T}$ by $\Gamma_{i}$, since both sets can be chosen as orthonormal set for the space of Hermitian traceless matrices on $H_{d_{1}}$. Given a general bi-partite density matrix as
$\rho=\frac{1}{d_{1}d_{2}}I_{d_{1}}\otimes I_{d_{2}}+{a}_{i}\Gamma_{i}\otimes I_{d_{2}}+{b}_{\mu}I_{d_{1}}\otimes\Omega_{\mu}+{r}_{i\mu}\Gamma_{i}\otimes\Omega_{\mu},$ (25)
one finds that
$\operatorname{Tr}(W\rho)=1+d_{1}{\bf b}\cdot{\bf t}+d_{1}\operatorname{Tr}(R\Lambda).$ (26)
@@
Those EW’s which correspond to unital positive maps (i.e. ${\bf t}=0$) are called unital entanglement witness. Therefore for any two dimensions, any operator of the above form can be an entanglement witness, provided that the condition (12) holds. Whether or not an EW constructed in this way, has also a negative eigenvalue depends on the matrix $\Lambda$ that we choose. In all previous papers that were mentioned in the introduction, $\Lambda$ should be orthogonal. These comprise a much smaller class compared to the $\Lambda$ matrices in this paper. Moreover, finding orthogonal matrices in high dimensions is not a straightforward task. In this paper, one can use any matrix with any properties to construct entanglement witnesses, as long as eq. (12) is satisfied. For any given density matrix $\rho$ characterized by the pair $({\bf b},R)$ as shown above, we have to find a witness characterized by a pair $({\bf t},\Lambda)$ such that $\operatorname{Tr}(W\rho)$ as written above is minimized. Note that for this kind of witness the parameter ${\bf a}$ in the density matrix plays no role. This minimization should be done in the range of all possible entanglement witnesses, namely among all those affine transormations which satisfy the inequality (65), which we rewrite here for convenience:
$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}}.$ (27)
To proceed, we need the following:
Definiiton: A unital entanglement witness is one which corresponds to the Choi matrix of a unital positive map.
We now show that show that for any given state $\rho$, if a non-unital EW can detect its entanglement, there is a unital EW whcih can detect it in a more effective way. This provides a rigorous justification for the focus of researchers in the past years on trying to construct unital positive maps and their corresponding entanglement witnesses.
Theorem: Given a general bi-partites state $\rho$, unital entanglement witnesses (e.g. those derived from unital maps) are more effective in detecting entanglement of this state.
Proof: In $\operatorname{Tr}(W\rho)$ given in (26) the vector ${\bf b}$ and the matrix $R$ are determined by the density matrix $\rho$ and our task is to choose the vector ${\bf t}$ and the matrix $\Lambda$ to make $\operatorname{Tr}(W\rho)$ as negative as possible. To do so, we invoke the Chauchy Schwartz inequality and choose ${\bf t}$ anti-parallel to ${\bf b}$, that is we take ${\bf t}=-\alpha{\bf b}$, where $\alpha$ is positive constant which can be taken as large as possible provided that the condition (12), namely
$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}}$
is not violated.
To minimize $\operatorname{Tr}(\Lambda R)$, let us use the singular value decomposition of $R$, i.e. $R=USV$, where $U,V$ are rotation matrices respectively belonging to orthogonal groups $O(d_{1})$ and $O(d_{2})$ and $S\in M_{d_{1}\times d_{2}}$ is a rectangular matrix on whose main diagonal are the singular values of $R$. Then using the freedom of choosing $\Lambda$, we write it as $\Lambda=V^{T}LU^{T}$, where we should choose $L$ such that $\operatorname{Tr}(\Lambda R)$ is minimized subject to the condition (12). Thus we have
$\operatorname{Tr}(\Lambda R)=\operatorname{Tr}(LS).$ (28)
Noting that the singular values of $R$ (i.e. the diagonal entries of $S$) are all positive, this equation can be minimized by considering $L$ to have on its diagonal, a non-negative parameter denoted by $-\beta$ and all the other off-diagonal elements to be zero. As explained in detail in appendix I, this leads to $\operatorname{Tr}(R\Lambda)=-\beta\|R\|_{1}.$ We then find
$\operatorname{Tr}(W\rho)=1-\alpha\|{\bf b}\|^{2}-d_{1}\beta\|R\|_{1}.$ (29)
We have to find the minimum of this quantity in the region $\alpha-\beta$ defined by
$\alpha\|{\bf b}\|+\sqrt{d_{1}(d_{1}-1)}\beta\leq\sqrt{\frac{d_{2}}{d_{2}-1}}.$ (30)
This is now a simple mathematical problem which can be expressed most simply when we cast it as follows: We want to find the minimum value of the function $F(\alpha,\beta):=\operatorname{Tr}(\rho W)$
$F(\alpha,\beta):=1-p\alpha-q\beta,$
in the region
$p^{\prime}\alpha+q^{\prime}\beta\leq 1,$
where
$p=\|{\bf b}\|^{2},\ \ \ q=d_{1}\|R\|_{1},\ \ \ p^{\prime}={\|{\bf b}\|}\sqrt{\frac{d_{2}-1}{d_{2}}},\ \ \ q^{\prime}=\sqrt{\frac{d_{1}(d_{1}-1)(d_{2}-1)}{d_{2}}}.$ (31)
The function $F$ being a linear function of $\alpha$ and $\beta$ aquires its minimum on one of the three vertices of the triangle, figure (2).
Figure 2: We should minimize the amount of $\operatorname{Tr}(W\rho)=1-p\alpha-q\beta$ with respect to the gray area, where $p,q,p^{\prime},q^{\prime}$ are difined in eq. (31).
On the vertex $(0,0)$, we have $F(0,0)=1,$ which is positive, hence no detection of entanglement. To explore the value of $F$ on the other two vertices, we calculate the value of $F$ on the oblique edge of the triangle defined by $\alpha p^{\prime}+\beta q^{\prime}=1$. On this edge, the function $F$ depends on one parameter $\alpha$, which after eliminating $\beta$ in favor of $\alpha$, allows us to define the following one-parameter function $\tilde{F}(\alpha)=\operatorname{Tr}(\rho W)$
$\tilde{F}(\alpha)=1-p\alpha-q(\frac{1-\alpha p^{\prime}}{q^{\prime}})=1-\frac{q}{q^{\prime}}-\alpha p^{\prime}(\frac{p}{p^{\prime}}-\frac{q}{q^{\prime}}).$ (32)
Note that the parameters $p,p^{\prime},q$ and $q^{\prime}$ come from the density matrix $\rho$ and $\alpha$ comes from the witness $W$ which is to be chosen so that the value of $\tilde{F}$ is as negative as possible. In case that $\frac{p}{p^{\prime}}<\frac{q}{q^{\prime}}$, we take $\alpha$ to be zero leading to the minimum value
$\tilde{F}_{min}=\tilde{F}(\alpha=0)=1-\frac{q}{q^{\prime}}.$ (33)
In case that $\frac{p}{p^{\prime}}>\frac{q}{q^{\prime}}$, we take $\alpha$ to be as large as possible, namely $\frac{1}{p^{\prime}}$ which gives the minimum value
$\tilde{F}_{min}=\tilde{F}(\alpha=\frac{1}{p^{\prime}})=1-\frac{p}{p^{\prime}}.$ (34)
But the condition $\frac{p}{p^{\prime}}<\frac{q}{q^{\prime}}$ already shows that $\tilde{F}(\alpha=0)\leq\tilde{F}(\alpha=\frac{1}{p^{\prime}})$ which means that there is no advantage in using non-unital witnesses, that is, any entangled state which can be detected by a non-unital witness ($\alpha=\frac{1}{q^{\prime}}$), is also detected by a unital witness ($\alpha=0$). Those states which are detected by this unital witness are those for which $\tilde{F}(0)=1-\frac{q^{\prime}}{q}<0$. Inserting the values of $q^{\prime}$ and $q$ from (31), this leads to the following condition:
${\rm If\ \ }\|R\|_{1}>\sqrt{\frac{(d_{1}-1)(d_{2}-1)}{d_{1}d_{2}}},\ \ \ \ \ \ {\rm the\ state\ is\ entangled.}$ (35)
@@
We stress the significance of this result which is now independent of any entanglemet witnesses and is a criterion based entirely on the density matrix itself. We will now show by several examples that this criterion, originated from the EW defined in (24), is indeed powerful enough to detect entanglement of a class of PPT entangled states and indeed in some examples, the range of parameters where the state is shown to be entangled is much larger than previously reported.
@@ 5 Examples
In this section, we consider a few examples and show the power of our criterion. This criterion which is derived here by starting from Mehta’s lemma and positive maps, is a special case of a more general criteria already reported in [21]. In all the following examples, except example 5.1, the states are PPT entangled states reported in previous works [22, 23]. We will show that the criterion given in (35), can detect entanglement in a larger interval of the parameters previously reported, and it also can detect PPT entangled states with different dimensions, i.e. $d_{1}\neq d_{2}$. For checking the following examples, the reader should note that we have used the following order for Gellmann matrices:
@@ 5.1 Example 1:
It is readily found that
@@
$\|R\|_{1}=\operatorname{Tr}(\sqrt{R^{T}R})=\frac{p}{d}(d^{2}-1),$
which, according to eq (35), should be greater than $\frac{d-1}{d}$ for the state to be entangled. Thus the range of $p$ for the state to be entangled is $p>\frac{1}{d+1}$. This is the same value obtained from the Peres criterion which in this case follows from